# A Survey on LoRA of Large Language Models

### v1

@@ 4.3 Mixture of LoRA Experts
When the LoRA plugins are trainable, we can jointly learn the mixture weights and the LoRA plugins, which can further improve the performance of the LoRA mixture. To jointly learn the mixture weights and LoRA plugins, Mixture of LoRA Experts (LoRA MoE) is a natural choice, where each LoRA plugin acts as an expert, while a router network typically assigns the mixture weights. LoRA MoE has been proven to be effective in many tasks, such as continual learning[63, 64], vision-language tasks[65] and multi-task medical applications[66].
@@ 5 Efficiency Improving
With the popularization of LLMs, the demand for training and running LoRA plugins increases rapidly. This increasing demand brings an unignorable computational burden; thus, for LoRA, the smaller, the faster, the better. To meet this demand, existing methods improve the computational efficiency of LoRA from the perspectives of (1) parameter reduction; (2) parameter quantization; (3) parallel LoRA computing frameworks.
@@ 5.1.3 Parameter Sharing
Parameter-sharing methods reduce the number of parameters by sharing parameters across different layers or modules of LLMs. VeRA[86] and VB-LoRA[87] are two representative parameter-sharing methods for LoRA. Specifically, VeRA proposes to share a pair of frozen random matrices across all layers and conduct layer-wise adaptation with “scaling vectors”. By contrast, VB-LoRA proposes a “divide-and-share” paradigm, which divides LoRA’s low-rank decomposition by a rank-one decomposition and achieves global sharing based on an admixture model.
@@ 5.2.1 PTQ-based methods
In PTQ-based methods, we first quantize an LLM and then fine-tune the quantized model, namely quantization and fine-tuning are sequentially conducted. QLoRA [88] is the first PTQ-based quantization-aware LoRA method. In the fine-tuning stage, it first quantizes an LLM to 4 bits and then fine-tunes a LoRA plugin on it with a higher precision, such as BFloat16 or Float16. In the inference stage, it dequantizes the LLM to the same precision as LoRA and then adds the LoRA updates to the LLM.
Although QLoRA can significantly reduce memory cost for fine-tuning, it does not bring benefits for inference, because it requires dequantizing the LLM to high precision again. To solve this problem, QA-LoRA [89] is proposed to reduce memory cost for both the fine-tuning and inference stages. QA-LoRA uses group-wise operators to balance the degrees of freedom of the LLM quantization and fine-tuning, which enables it to obtain a LoRA plugin having identical precision with the quantized LLM. Thus, it can perform inference without dequantization.
@@ 5.3 Parallel LoRA Computing Frameworks
LoRA’s parameter-efficient nature enables us to fine-tune or infer multiple plugins on a single GPU or a GPU cluster, which can save computational resources and improve the efficiency of LoRA. This section introduces the parallel fine-tuning and parallel inference frameworks, respectively.
@@ 5.3.1 Parallel Fine-tuning
Parallelly fine-tuning multiple LoRA plugins on a single GPU can reduce GPU memory usage and improve computation efficiency. ASPEN [93] proposes a high-throughput parallel finetuning framework for LoRA, which consists of a BatchFusion approach and an adaptive job scheduling algorithm. Specifically, the BatchFusion approach supports parallelly fine-tuning multiple LoRA plugins on a shared LLM by fusing multiple input batches into a single batch, while the adaptive job scheduling algorithm allocates computation resources to the fine-tuning jobs.
@@ 5.3.2 Parallel Inference
Parallel inference framework for LoRA can not only improve the computational efficiency but also support the needs of multi-tenant service. Punica[94] uses a new CUDA kernel design to batch GPU operations for different LoRA plugins. Based on Punica, S-LoRA [95] further optimizes the parallel inference framework by introducing a unified paging mechanism and a new tensor parallelism strategy, which enables the service of thousands of concurrent LoRA plugins. Then, based on Punica and S-LoRA, CARASERVE [96] reduces the cold-start overhead and further improves the service efficiency and SLO (service-level objective) attainment rates by CPU-GPU cooperation and rank-aware scheduling.
@@ 8 Conclusion and Future Direction
In this survey, the recent progress of LoRA have been systematically reviewed from the perspective of downstream adaptation improving, cross-task generalization, efficiency improving, federated learning and applications. From this review, we can find that LoRA is parameter efficient, pluggable, compatible and easy to achieve cross-task generalization, which enables it to be one of the most important technology for LLMs applications. Recent progress further boost the generalization and efficiency of LoRA, and stimulate its potential to be used in more scenarios. Here, we list three future directions where LoRA will be indispensable.
@@ 8.1 LoRA for GaaS
In Generative-as-a-Service (GaaS), cloud-based platforms provide users with generative artificial intelligence (AGI) services. GaaS enables users enjoy AGI without deploying local computational resources. For the users’ needs are diverse, it is necessary to provides various functions for GaaS. To implement the various functions, we can construct a LoRA plugin for each function. The pramameter efficiency and plugability of LoRA can facilitate efficient functions’ construction and execution. Besides, the services on GaaS platforms can change rapidly alonging time. To follow the changes, we can train new LoRA plugins that initialized by combination of previous plugins. The cross-task generalization ability of LoRA can facilitate fast adaption to service updations.
@@ 8.3 LoRA for Autonomous Agents
In the current agents, the roles are typically assigned by prompts; however, prompt may cannot give a comprehensive discription of the role when the role is complex and the number of related data is large. Assiging roles with LoRA plugins training from data related to the roles can be a better choice. Furthermore, the tools for agent can be LoRA plugins. Besides, the memory usually augments the agents with retrieval augmented generation (RAG); however, due to the input token limitation and the short-comings of in-context learning, the RAG-based support may be less effective. By contrast, we can use LoRA-based continual learning to construct memory plugins, which can solve the problem of RAG. Therefore, LoRA-driven agents are worth to explore.

### v2

@@ 4.3 Mixture of LoRA Experts
When the LoRA modules are trainable, we can jointly learn the mixture weights and the LoRA modules, which can further improve the performance of the LoRA mixture. To jointly learn the mixture weights and LoRA modules, Mixture of LoRA Experts (LoRA MoE) is a natural choice, where each LoRA module acts as an expert, while a router network typically assigns the mixture weights. LoRA MoE has been proven to be effective in many tasks, such as continual learning[65, 66], vision-language tasks[67] and multi-task medical applications[68].
@@ 5 Efficiency Improving
With the popularization of LLMs, the demand for training and running LoRA modules increases rapidly. This increasing demand brings an unignorable computational burden; thus, for LoRA, the smaller, the faster, the better. To meet this demand, existing methods improve the computational efficiency of LoRA from the perspectives of (1) parameter reduction; (2) parameter quantization; (3) parallel LoRA computing frameworks.
@@ 5.1.3 Parameter Sharing
Parameter-sharing methods reduce the number of parameters by sharing parameters across different layers or modules of LLMs. VeRA[88] and VB-LoRA[89] are two representative parameter-sharing methods for LoRA. Specifically, VeRA proposes to share a pair of frozen random matrices across all layers and conduct layer-wise adaptation with “scaling vectors”. By contrast, VB-LoRA proposes a “divide-and-share” paradigm, which divides LoRA’s low-rank decomposition by a rank-one decomposition and achieves global sharing based on an admixture model. Instead of sharing parameters in the original parameter space, FourierFT [90] converts the incremental matrix $\Delta{W}$ into the spatial domain using Fourier transform. It shares spectral entries across all layers and only learns its sparse spectral coefficients for each layer, thus reducing the number of trainable parameters.
@@ 5.2.1 PTQ-based methods
In PTQ-based methods, we first quantize an LLM and then fine-tune the quantized model, namely quantization and fine-tuning are sequentially conducted. QLoRA [91] is the first PTQ-based quantization-aware LoRA method. In the fine-tuning stage, it first quantizes an LLM to 4 bits and then fine-tunes a LoRA module on it with a higher precision, such as BFloat16 or Float16. In the inference stage, it dequantizes the LLM to the same precision as LoRA and then adds the LoRA updates to the LLM.
Although QLoRA can significantly reduce memory cost for fine-tuning, it does not bring benefits for inference, because it requires dequantizing the LLM to high precision again. To solve this problem, QA-LoRA [92] is proposed to reduce memory cost for both the fine-tuning and inference stages. QA-LoRA uses group-wise operators to balance the degrees of freedom of the LLM quantization and fine-tuning, which enables it to obtain a LoRA module having identical precision with the quantized LLM. Thus, it can perform inference without dequantization.
@@ 5.3 Parallel LoRA Computing Frameworks
LoRA’s parameter-efficient nature enables us to fine-tune or infer multiple modules on a single GPU or a GPU cluster, which can save computational resources and improve the efficiency of LoRA. This section introduces the parallel fine-tuning and parallel inference frameworks, respectively.
@@ 5.3.1 Parallel Fine-tuning
Parallelly fine-tuning multiple LoRA modules on a single GPU can reduce GPU memory usage and improve computation efficiency. ASPEN [96] proposes a high-throughput parallel finetuning framework for LoRA, which consists of a BatchFusion approach and an adaptive job scheduling algorithm. Specifically, the BatchFusion approach supports parallelly fine-tuning multiple LoRA modules on a shared LLM by fusing multiple input batches into a single batch, while the adaptive job scheduling algorithm allocates computation resources to the fine-tuning jobs.
@@ 5.3.2 Parallel Inference
Parallel inference framework for LoRA can not only improve the computational efficiency but also support the needs of multi-tenant service. Punica[97] uses a new CUDA kernel design to batch GPU operations for different LoRA modules. Based on Punica, S-LoRA [98] further optimizes the parallel inference framework by introducing a unified paging mechanism and a new tensor parallelism strategy, which enables the service of thousands of concurrent LoRA modules. Then, based on Punica and S-LoRA, CARASERVE [99] reduces the cold-start overhead and further improves the service efficiency and SLO (service-level objective) attainment rates by CPU-GPU cooperation and rank-aware scheduling.
@@ 6.4 Parameter Privacy
Furthermore, recent works have integrated differential privacy (DP) techniques with LoRA in federated learning to enhance data privacy. DP-LoRA [218] ensures differential privacy by adding Gaussian noise to LoRA’s weight updates during the update process. This approach maintains privacy and improves communication efficiency. To solve the noise amplification when applying differential privacy in LoRA, FFA-LoRA [102] fixes the matrix $A$, avoiding the local semi-quadratic structure and enhancing robustness and performance.
@@ 8 Conclusion and Future Direction
In this survey, the recent progress of LoRA have been systematically reviewed from the perspective of downstream adaptation improving, cross-task generalization, efficiency improving, federated learning and applications. From this review, we can find that LoRA is parameter efficient, pluggable, compatible and easy to achieve cross-task generalization, which enables it to be one of the most important technology for LLMs applications. Recent progress further boosts the generalization and efficiency of LoRA, and stimulate its potential to be used in more scenarios. Here, we list three future directions where LoRA will be indispensable.
@@ 8.1 LoRA for GaaS
In Generative-as-a-Service (GaaS), cloud-based platforms provide users with generative artificial intelligence (AGI) services. GaaS enables users enjoy AGI without deploying local computational resources. For the users’ needs are diverse, it is necessary to provides various functions for GaaS. To implement the various functions, we can construct a LoRA module for each function. The pramameter efficiency and plugability of LoRA can facilitate efficient functions’ construction and execution. Besides, the services on GaaS platforms can change rapidly alonging time. To follow the changes, we can train new LoRA modules that initialized by combination of previous LoRA modules. The cross-task generalization ability of LoRA can facilitate fast adaption to service updations.
@@ 8.3 LoRA for Autonomous Agents
In the current agents, the roles are typically assigned by prompts; however, prompt may cannot give a comprehensive discription of the role when the role is complex and the number of related data is large. Assiging roles with LoRA modules training from data related to the roles can be a better choice. Furthermore, the tools for agent can be LoRA modules. Besides, the memory usually augments the agents with retrieval augmented generation (RAG); however, due to the input token limitation and the short-comings of in-context learning, the RAG-based support may be less effective. By contrast, we can use LoRA-based continual learning to construct memory modules, which can solve the problem of RAG. Therefore, LoRA-driven agents are worth to explore.